OpenStudy (anonymous):

What is the definition of a equation And give an example with a expression shown

OpenStudy (igloo3000):

An equation is when you have an equal sign like 8+2=10 or 5+5=10

OpenStudy (igloo3000):

it is also equal to something

OpenStudy (anonymous):

an equation is identify by an equal sign. .(=)

OpenStudy (anonymous):

How about an example "jerwyn gayo".

OpenStudy (igloo3000):

8-5=3